OpenStudy (anonymous):

say that a and b are the lengths of the legs of a right triangle. and c is the length os the hypotenuse. find c if a = 12 and b =16 a. 40 b. 10 c. 400 d. 20

OpenStudy (anonymous):

12^2+16^2=c^2 144+256=400 square root of 400 = 20 c=20